# Drift2Matrix: Kernel-Induced Self Representation for Concept Drift Adaptation in Co-evolving Time Series

### v2

@@ 6.5 Additional Experiments
Further experiments and ablation studies can be found in Appendix H, including
  • Appendix H.1: We evaluate Drift2Matrix’s ability to handle noise or outliers, demonstrating its robustness.
  • Appendix H.2: We expanded the model’s evaluation to include the other datasets like MSP, ELD, CCD, EQD, EOG, and RDS. For each dataset, distinct concepts exhibited by co-evolving series were identified, demonstrating the model’s robustness in concept identification across various datasets.
  • Appendix H.3: A complete result comparing our model with other models, providing a comprehensive view of the model’s performance across all datasets.
  • Appendix H.4: A comparative analysis was conducted between our model and the N-BEATS model on Stock1 and Stock2 datasets. This comparison highlighted the limitations of N-BEATS in capturing complex concept transitions within multiple time series.
  • Appendix H.5: The model’s application to motion segmentation on the Hopkins155 database was explored. This case study demonstrated the model’s effectiveness in handling different types of sequences and its adaptability to various motion concepts.
  • Appendix H.6: A comprehensive analysis of RMSE values across all datasets was presented, showcasing our model’s superior performance in comparison to other models.
  • Appendix H.7: A comprehensive analysis of complexity and execution time Evaluation.
  • Appendix H.8: Detailed ablation studies were conducted to validate the efficacy of various components of the Drift2Matrix’s kernel representation learning, including regularizations, kernel-based methods, and different kernel functions. These studies provided insights into the model’s performance under different configurations and conditions.
@@ 7 Conclusion
Despite its strengths, Drift2Matrix has a limitation when applied to time series with few variables. For example, converting a dataset with five variables into a 5x5 matrix makes block diagonal regularization less effective. Conversely, larger datasets, like those with 500 variables, benefit significantly from our method, enabling the identification of nonlinear relationships and concept drift. This characteristic is somewhat counterintuitive compared to most existing time series models that often focus on single or low-dimensional (few variables) time series forecasting, such as sensor data streams. Despite this limitation, we believe it underscores Drift2Matrix’s unique appeal. It addresses a gap in handling concept drift in time series with a large number of variables, offering excellent interpretability and reduced computational complexity.
Supplementary Materials for: Drift2Matrix
@@ Appendix A Extended Related Work and Motivation
Concept drift models. Co-evolving time series analysis, by its nature, entails the simultaneous observation and interpretation of interdependent data streams Cavalcante et al. (2016). This complexity is further heightened when concept drift is introduced into the model Webb et al. (2016). In such scenarios, a shift in one variable can propagate through the network of interrelations, affecting the entire co-evolving system. Matsubara et al. Matsubara & Sakurai (2016) put forth the RegimeCast model, which learns potential patterns within a designated time interval in a co-evolving environment and predicts the subsequent pattern most likely to emerge. While the approach can forecast following patterns, it is not designed to account for any interdependencies between them. In their subsequent work Matsubara & Sakurai (2019), the authors introduced the deterministic OrbitMap model to capture the temporal transitions across displayed concepts. Notably, this approach relies on pre-labeled concepts (known beforehand). DDG-DA Li et al. (2022) for data distribution generation has been adapted to better suit co-evolving scenarios, addressing the unique challenges presented by the interplay of multiple data streams under concept drift conditions. However, this method defines the concept as a collective behavior represented by co-evolving time series, rather than capturing the dynamics of individual series and their interactions. While acknowledging that deep learning has made significant advances in time series field, we must also note that most of these progress aims at improving accuracy. For example, OneNet Wen et al. (2024) addresses the concept drift problem by integrating an ensemble of models that share different data biases and learning to dynamically combine forecasts from these models for enhanced prediction. It maintains two forecasting models focusing on temporal correlation and cross-variable dependency, trained independently and dynamically adjusted during testing; FSNet Pham et al. (2022), on the other hand, is designed to quickly adapt to new or recurring patterns in non-stationary environments by enhancing a neural network backbone with two key components: an adapter for recent changes and an associative memory for recurrent patterns. Dish-TS Fan et al. (2023) offers a general approach for alleviating distribution shift in time series forecasting by normalizing model inputs and outputs to better handle distribution changes. Similarly, Cogra’s application of the Sequential Mean Tracker (SMT) adjusts to changes in data distribution, improving forecast accuracy Miyaguchi & Kajino (2019).
@@
Representation Learning on TS. Representation learning on time series (TS) has gained significant attention due to its potential in uncovering underlying patterns and features essential for various downstream tasks. T-Rep Fraikin et al. (2023) leverages time-embeddings for time series representation. This method focuses on capturing temporal dependencies and variations through time-specific embeddings. TimesURL Liu & Chen (2024) employs self-supervised contrastive learning to create representations for time series. BTSF Yang & Hong (2022) is an unsupervised method for time series representation learning that iteratively fuses temporal and spectral information. The bilinear fusion mechanism allows the model to capture both temporal dynamics and spectral characteristics of the time series. Timemae Cheng et al. (2023) leverages self-supervised learning to learn representations of time series using decoupled masked autoencoders. This method focuses on reconstructing the time series data by masking certain parts of the input and learning to predict the missing information. While the aforementioned methods have advanced time series representation learning, they have several limitations. Many approaches assume linear relationships, limiting their ability to capture complex, non-linear dependencies inherent in co-evolving time series. Additionally, techniques heavily reliant on specific features, such as spectral characteristics, may not generalize well across diverse datasets. The computational complexity of some advanced representation learning methods poses challenges for scalability, especially when applied to large, co-evolving datasets. Moreover, these methods often focus on representation learning for single time series or treat co-evolving time series as a data stream, rather than uncovering the intricate non-linear relationships among series.
@@
Motivation. Our work aims to propose a novel perspective on concept evaluation in co-evolving time series. Eschewing traditional methods that rely on latent variable dynamics, we delve into the inherent behavior of the time series. Our proposed Drift2Matrix, with its nonlinear mapping, is adept at capturing the ever-changing concepts, offering insights into their intricacies and forecasting potential concept drift. The comparison of our Drift2Matrix framework with the recent advances in deep learning methods, is given from the following three perspectives:
@@
  • Focus of Research. While acknowledging the rapid developments in deep learning for time series forecasting, it’s crucial to point out that most of these advancements concentrate predominantly on forecasting accuracy. Notably, even the most recent deep learning methods that mention concept drift, such as OneNet Wen et al. (2024) and FSNet Pham et al. (2022), primarily aim to mitigate the impact of concept drift on forecasting. They achieve this by incorporating an ensemble of models with diverse data biases or by refining network parameters for better adaptability. In contrast, Drift2Matrix focuses on the challenges of adaptive concept identification and dynamic concept drift in co-evolving time series. Our model delves deeper into the inherent structure of time series data, enabling a more nuanced understanding and handling of concept drift by dynamically identifying and adapting to new concepts as they emerge.
@@
  • Interpretability. Drift2Matrix introduces kernel-induced representation to reveal nonlinear relationships in time series, substantially boosting both adaptability and interpretability of the model. In particular, Drift2Matrix transforms time series data into a matrix format, where its block diagonal structure intuitively maps out distinct concepts. Conversely, while methods like feature importance scoring and attention mechanisms aim to improve deep learning models’ interpretability, they often rely on post-hoc analysis of the model’s internal mechanisms.
@@
  • Example of Financial Market. Financial time series analysis transcends mere forecasting; it demands interpretability that builds trust and supports applications like portfolio management. Drift2Matrix excels in this regard, offering clear insights into market dynamics beyond the conventional categories of bull, bear, or sideways markets. It adeptly captures a wide array of market scenarios, identifying distinct concepts driven by various factors—whether it’s value versus growth or the interplay between small and large caps. For in-depth case studies, including analyses of the Stock1 and Stock2 datasets, please refer to Sec 6.2 & 6.4.
@@
Formal Mathematical Definition of Concepts. For window $W_{p}$, the $r$-th concept is defined as the vector representation of subseries corresponding to the $r$-th block, $\mathbf{Z}_{p}^{(r)}$, in the representation matrix $\mathbf{Z}_{p}$. Specifically, it aligns with the centroid of similar subseries, represented as $C_{r,p}=\text{Centroid}\Big{(}\{\mathbf{S}_{i}\ |\ \mathbf{S}_{i}\in\mathbf{Z}_{p}^{(r)}\}\Big{)}$.To differentiate and refine similar or repeated concepts across different windows, two concepts $C_{r,p}$ and $C_{s,p+1}$ are considered distinct if $||C_{r,p}-C_{s,p+1}||\_F^{2}>\rho$, where $\rho$ is a tunable hyperparameter that regulates the granularity of concept.
@@ Appendix B Estimating the Number of Concepts
In time series analysis, accurately identifying the number of concepts, such as periods of varying volatility in financial market, stages of a disease in medical monitoring, or climatic patterns in meteorology, is crucial. These concepts offer insights for data-driven decision-making, understanding underlying dynamics, and predicting future behaviors. While estimating the number of concepts is generally challenging, our kernel-induced representation learning approach offers a promising solution. Leveraging the block-diagonal structure of the self-representation matrix produced by our method, we can effectively estimate the number of concepts. According to the Laplacian matrix property Von Luxburg (2007), a strictly block-diagonal matrix $\mathbf{Z}$ allows us to determine the number of concepts $k$ by first calculating the Laplacian matrix of $\mathbf{Z}$ ($\mathbf{L_{Z}}$) and then counting the number of zero eigenvalues of $\mathbf{L_{Z}}$. Although the dataset is not always clean or noise-free (as is often the case in practice), we propose an eigengap thresholding approach to estimating the number of concepts. This approach estimates the number of concepts $\hat{k}$ as:
@@
Here, ${\lambda_{i}}_{i=1}^{N}$ are the eigenvalues of $\mathbf{L_{Z}}$ in increasing order. The eigengap, or the difference between the $i^{th}$ and $(i+1)^{th}$ eigenvalues, plays a crucial role. According to matrix perturbation theory Stewart (1990), a larger eigengap indicates a more stable subspace composed of the selected $k$ eigenvectors. Thus, the number of concepts can be determined by identifying the first extreme value of the eigengap  [^5] In this paper, we simply initialize the number of concepts for each window as 3 to obtain the initial representation $\mathbf{Z}$, and then estimate $k$..
@@ Appendix C Proofs and Optimization
C.1 Permutation Invariance of Representation Matrix in SeqMatrix
Theorem C.1
In Drift2Matrix, the representation matrix obtained for a permuted input data is equivalent to the permutation-transformed original representation matrix. Specifically, let $\mathbf{Z}$ be feasible to $\Phi(\mathbf{S})=\Phi(\mathbf{S})\mathbf{Z}$, then $\tilde{\mathbf{Z}}=\mathbf{P^{T}ZP}$ is feasible to $\Phi(\tilde{\mathbf{S}})=\Phi(\tilde{\mathbf{S}})\tilde{\mathbf{Z}}$.
Proof C.2
Given a permutation matrix $\mathbf{P}$, consider the self-representation matrix $\tilde{\mathbf{Z}}$ for the permuted data matrix $\mathbf{S}\mathbf{P}$. The objective for $\mathbf{S}\mathbf{P}$ becomes:
$\begin{split}\min_{\tilde{\mathbf{Z}}}\left\|\Phi(\mathbf{SP})-\Phi(\mathbf{SP})\tilde{\mathbf{Z}}\right\|^{2}+\Omega(\tilde{\mathbf{Z}}),\quad\text{s.t.}\ \tilde{\mathbf{Z}}=\tilde{\mathbf{Z}}^{\mathrm{T}}\geq 0,\ \text{diag}(\tilde{\mathbf{Z}})=0\end{split}$ (10)
By the properties of kernel functions and permutation matrices, we have $\Phi(\mathbf{SP})=\Phi(\mathbf{S}P)$. Substituting this into the objective function for $\tilde{\mathbf{Z}}$, we have:
$\begin{split}\min_{\tilde{\mathbf{Z}}}\left\|\Phi(\mathbf{S}\mathbf{P})-\Phi(\mathbf{S})\mathbf{P}\tilde{\mathbf{Z}}\right\|^{2}+\Omega(\tilde{\mathbf{Z}})\quad\text{s.t.}\ \tilde{\mathbf{Z}}=\tilde{\mathbf{Z}}^{\mathrm{T}}\geq 0,\ \text{diag}(\tilde{\mathbf{Z}})=0\end{split}$ (11)
Since $P$ is a permutation matrix,$\mathbf{P}\mathbf{P}^{\mathrm{T}}=\mathbf{I}$, the identity matrix. We apply the transformation $\mathbf{P\tilde{Z}P^{\mathrm{T}}}$ to the objective function:
$\begin{split}\min_{\tilde{\mathbf{Z}}}\left\|\Phi(\mathbf{S})-\Phi(\mathbf{S})\mathbf{P\tilde{Z}P^{\mathrm{T}}}\right\|^{2}+\Omega(\tilde{\mathbf{Z}})\quad\text{s.t.}\ \tilde{\mathbf{Z}}=\tilde{\mathbf{Z}}^{\mathrm{T}}\geq 0,\ \text{diag}(\tilde{\mathbf{Z}})=0\end{split}$ (12)
For the function to be minimized, $\mathbf{P\tilde{Z}P^{\mathrm{T}}}$ must be the optimal representation matrix for $\mathbf{S}$, which is $\mathbf{Z}$. Therefore, $\mathbf{P\tilde{Z}P^{\mathrm{T}}}=\mathbf{Z}$, or equivalently, $\tilde{\mathbf{Z}}=\mathbf{P}^{\mathrm{T}}\mathbf{ZP}$.
This result shows that the self-representation matrix $\mathbf{Z}$ for $\mathbf{S}$ transforms to $\mathbf{P}^{\mathrm{T}}\mathbf{ZP}$ for the permuted data matrix $\mathbf{SP}$, demonstrating the invariance of the representation matrix under permutations of the data matrix.
C.2 Manifold Structure Preservation in Drift2Matrix
Theorem C.3
Drift2Matrix reveals nonlinear relationships among time series in a high-dimensional space while simultaneously preserving the local manifold structure of series.
Proof C.4
Optimization problem Eq. 3 can be converted to the form of a matrix trace:
$\min_{\mathbf{Z}}\frac{1}{2}\mathrm{Tr}(\bm{\mathcal{K}}-2\bm{\mathcal{K}}\mathbf{Z}+\mathbf{Z}^{\mathrm{T}}\bm{\mathcal{K}}\mathbf{Z})+\gamma||\mathbf{Z}||_{\boxed{\scriptstyle k}},$ (13)
In the above, the negative term $-\mathrm{Tr}(\bm{\mathcal{K}}\mathbf{Z})$ can be transformed into
$\min_{\mathbf{Z}}-\mathrm{Tr}(\bm{\mathcal{K}}\mathbf{Z})$ $=\min_{\mathbf{Z}}\sum_{i=1}^{N}\sum_{j=1}^{N}-\mathrm{\Phi}(\mathbf{S}_{i})^{\mathrm{T}}\mathrm{\Phi}(\mathbf{S}_{j})\mathbf{Z}_{ij}=\min_{\mathbf{Z}}\sum_{i=1}^{N}\sum_{j=1}^{N}-\bm{\mathcal{K}}(\mathbf{S}_{i},\mathbf{S}_{j})\mathbf{Z}_{ij}$ (14)
where $\bm{\mathcal{K}}(\mathbf{S}_{i},\mathbf{S}_{j})$ indicates the similarity between $\mathbf{S}_{i}$ and $\mathbf{S}_{j}$ in kernel space. It can be seen from Eq. 14 that a large similarity (small distance) $\bm{\mathcal{K}}(\mathbf{S}_{i},\mathbf{S}_{j})$ tends to cause a large $\mathbf{Z}_{ij}$, and vice versa. This is in fact an kernel extension of preserving local manifold structure in linear space, i.e., $\min_{\mathbf{Z}}\sum_{i=1}^{N}\sum_{j=1}^{N}||\mathbf{X}_{i}-\mathbf{X}_{j}||^{2}\mathbf{Z}_{ij}$ Nie et al. (2014); Zhan et al. (2018). Suppose a small weight is given to this negative term, which means that the self-representation of the data will take into account the contribution of all other data. Conversely, it will only consider the contribution of other data that are nearest neighbours to the data, thus further enhancing the sparsity of the self-representation $\mathbf{Z}$ while maintaining the local manifold structure.
C.3 Optimization of Nonconvex Problem

### v3

@@ 6.5 Additional Experiments
Further experiments and ablation studies can be found in Appendix G, including
  • Appendix G.1: We evaluate CORAL’s ability to handle noise or outliers, demonstrating its robustness.
  • Appendix G.2: We expanded the model’s evaluation to include the other datasets like MSP, ELD, CCD, EQD, EOG, and RDS. For each dataset, distinct concepts exhibited by co-evolving series were identified, demonstrating the model’s robustness in concept identification across various datasets.
  • Appendix G.3: A complete result comparing our model with other models, providing a comprehensive view of the model’s performance across all datasets.
  • Appendix G.4: A comparative analysis was conducted between CORAL and the N-BEATS model on Stock1 and Stock2 datasets. This comparison highlighted the limitations of N-BEATS in capturing complex concept transitions within multiple time series.
  • Appendix G.5: The model’s application to motion segmentation on the Hopkins155 database was explored. This case study demonstrated the model’s effectiveness in handling different types of sequences and its adaptability to various motion concepts.
  • Appendix G.6: An analysis of RMSE values across all datasets was presented, showcasing our model’s superior performance in comparison to other models.
  • Appendix G.7: An analysis of complexity and execution time evaluation.
  • Appendix G.8: Detailed ablation studies were conducted to validate the efficacy of various components of the CORAL, including varying series number, regularizations, kernel-based methods, and different kernel functions. These studies provided insights into the model’s performance under different configurations and conditions.
@@ 7 Conclusion
Despite its strengths, CORAL has a limitation when applied to time series with few variables. For example, converting a dataset with five variables into a 5x5 matrix makes block diagonal regularization less effective. Conversely, larger datasets, like those with 500 variables, benefit significantly from our method, enabling the identification of nonlinear relationships and concept drift. This characteristic is somewhat counterintuitive compared to most existing time series models that often focus on single or low-dimensional (few variables) time series forecasting, such as sensor data streams. Despite this limitation, we believe it underscores CORAL’s unique appeal. It addresses a gap in handling concept drift in time series with a large number of variables, offering excellent interpretability and reduced computational complexity.
Impact Statements
This paper introduces a novel framework for modeling concept drift in co-evolving time series, aimed at advancing the field of Machine Learning. Our work enhances the adaptability and interpretability of time series models in dynamic environments. While it has potential applications in finance, healthcare, and climate modeling, we do not foresee any direct ethical concerns or societal risks associated with this research.
Supplementary Materials for: CORLA
@@ Appendix A Extended Related Work and Motivation
Concept drift models. Co-evolving time series analysis, by its nature, entails the simultaneous observation and interpretation of interdependent data streams (Cavalcante et al., 2016; Xu et al., 2024b, a). This complexity is further heightened when concept drift is introduced into the model (Webb et al., 2016). In such scenarios, a shift in one variable can propagate through the network of interrelations, affecting the entire co-evolving system. Matsubara et al. (Matsubara & Sakurai, 2016) put forth the RegimeCast model, which learns potential patterns within a designated time interval in a co-evolving environment and predicts the subsequent pattern most likely to emerge. While the approach can forecast following patterns, it is not designed to account for any interdependencies between them. In their subsequent work (Matsubara & Sakurai, 2019), the authors introduced the deterministic OrbitMap model to capture the temporal transitions across displayed concepts. Notably, this approach relies on pre-labeled concepts (known beforehand). DDG-DA (Li et al., 2022) for data distribution generation has been adapted to better suit co-evolving scenarios, addressing the unique challenges presented by the interplay of multiple data streams under concept drift conditions. However, this method defines the concept as a collective behavior represented by co-evolving time series, rather than capturing the dynamics of individual series and their interactions. While acknowledging that deep learning has made significant advances in time series field (Xu et al., 2024d, e, c), we must also note that most of these progress aims at improving accuracy. For example, OneNet (Wen et al., 2024) addresses the concept drift problem by integrating an ensemble of models that share different data biases and learning to dynamically combine forecasts from these models for enhanced prediction. It maintains two forecasting models focusing on temporal correlation and cross-variable dependency, trained independently and dynamically adjusted during testing; FSNet (Pham et al., 2022), on the other hand, is designed to quickly adapt to new or recurring patterns in non-stationary environments by enhancing a neural network backbone with two key components: an adapter for recent changes and an associative memory for recurrent patterns. Dish-TS (Fan et al., 2023) offers a general approach for alleviating distribution shift in time series forecasting by normalizing model inputs and outputs to better handle distribution changes. Similarly, Cogra’s application of the Sequential Mean Tracker (SMT) adjusts to changes in data distribution, improving forecast accuracy (Miyaguchi & Kajino, 2019).
@@
Representation Learning on TS. Representation learning on time series (TS) has gained significant attention due to its potential in uncovering underlying patterns and features essential for various downstream tasks. T-Rep (Fraikin et al., 2023) leverages time-embeddings for time series representation. This method focuses on capturing temporal dependencies and variations through time-specific embeddings. TimesURL (Liu & Chen, 2024) employs self-supervised contrastive learning to create representations for time series. BTSF (Yang & Hong, 2022) is an unsupervised method for time series representation learning that iteratively fuses temporal and spectral information. The bilinear fusion mechanism allows the
@@
model to capture both temporal dynamics and spectral characteristics of the time series. Timemae (Cheng et al., 2023) leverages self-supervised learning to learn representations of time series using decoupled masked autoencoders. This method focuses on reconstructing the time series data by masking certain parts of the input and learning to predict the missing information. While the aforementioned methods have advanced time series representation learning, they have several limitations. Many approaches assume linear relationships, limiting their ability to capture complex, non-linear dependencies inherent in co-evolving time series. Additionally, techniques heavily reliant on specific features, such as spectral characteristics, may not generalize well across diverse datasets. The computational complexity of some advanced representation learning methods poses challenges for scalability, especially when applied to large, co-evolving datasets. Moreover, these methods often focus on representation learning for single time series or treat co-evolving time series as a data stream, rather than uncovering the intricate non-linear relationships among series. Table 2 illustrates the relative advantages of our method.
@@
Motivation. Our work aims to propose a novel perspective on concept evaluation in co-evolving time series. Eschewing traditional methods that rely on latent variable dynamics, we delve into the inherent behavior of the time series. Our proposed CORAL, with its nonlinear mapping, is adept at capturing the ever-changing concepts, offering insights into their intricacies and forecasting potential concept drift. The comparison of our CORAL framework with the recent advances in deep learning methods, is given from the following three perspectives:
@@
  • Focus of Research. While acknowledging the rapid developments in deep learning for time series forecasting, it’s crucial to point out that most of these advancements concentrate predominantly on forecasting accuracy. Notably, even the most recent deep learning methods that mention concept drift, such as OneNet (Wen et al., 2024) and FSNet (Pham et al., 2022), primarily aim to mitigate the impact of concept drift on forecasting. They achieve this by incorporating an ensemble of models with diverse data biases or by refining network parameters for better adaptability. In contrast, CORAL focuses on the challenges of adaptive concept identification and dynamic concept drift in co-evolving time series. Our model delves deeper into the inherent structure of time series data, enabling a more nuanced understanding and handling of concept drift by dynamically identifying and adapting to new concepts as they emerge.
@@
  • Interpretability. CORAL introduces kernel-induced representation to reveal nonlinear relationships in time series, substantially boosting both adaptability and interpretability of the model. In particular, CORAL transforms time series data into a matrix format, where its block diagonal structure intuitively maps out distinct concepts. Conversely, while methods like feature importance scoring and attention mechanisms aim to improve deep learning models’ interpretability, they often rely on post-hoc analysis of the model’s internal mechanisms.
@@
  • Example of Financial Market. Financial time series analysis transcends mere forecasting; it demands interpretability that builds trust and supports applications like portfolio management (Chen et al., 2022a, b). CORAL excels in this regard, offering clear insights into market dynamics beyond the conventional categories of bull, bear, or sideways markets. It adeptly captures a wide array of market scenarios, identifying distinct concepts driven by various factors—whether it’s value versus growth or the interplay between small and large caps. For in-depth case studies, including analyses of the Stock1 and Stock2 datasets, please refer to Sec 6.2 & 6.4.
Formal Mathematical Definition of Concepts.
@@
For window $W_{p}$, the $r$-th concept is defined as the vector representation of subseries corresponding to the $r$-th block, $\mathbf{Z}_{p}^{(r)}$, in the representation matrix $\mathbf{Z}_{p}$. Specifically, it aligns with the centroid of similar subseries, represented as $C_{r,p}=\text{Centroid}\Big{(}\{\mathbf{S}_{i}\ |\ \mathbf{S}_{i}\in\mathbf{Z}_{p}^{(r)}\}\Big{)}$.To differentiate and refine similar or repeated concepts across different windows, two concepts $C_{r,p}$ and $C_{s,p+1}$ are considered distinct if $||C_{r,p}-C_{s,p+1}||_{F}^{2}>\rho$, where $\rho$ is a tunable hyperparameter that regulates the granularity of concept.
@@ Appendix B Estimating the Number of Concepts
In time series analysis, accurately identifying the number of concepts, such as periods of varying volatility in financial market, stages of a disease in medical monitoring, or climatic patterns in meteorology, is crucial. These concepts offer insights for data-driven decision-making, understanding underlying dynamics, and predicting future behaviors. While estimating the number of concepts is generally challenging, our kernel-induced representation learning approach offers a promising solution. Leveraging the block-diagonal structure of the self-representation matrix produced by our method, we can effectively estimate the number of concepts. According to the Laplacian matrix property (Von Luxburg, 2007), a strictly block-diagonal matrix $\mathbf{Z}$ allows us to determine the number of concepts $k$ by first calculating the Laplacian matrix of $\mathbf{Z}$ ($\mathbf{L_{Z}}$) and then counting the number of zero eigenvalues of $\mathbf{L_{Z}}$. Although the dataset is not always clean or noise-free (as is often the case in practice), we propose an eigengap thresholding approach to estimating the number of concepts. This approach estimates the number of concepts $\hat{k}$ as:
@@
Here, ${\lambda_{i}}_{i=1}^{N}$ are the eigenvalues of $\mathbf{L_{Z}}$ in increasing order. The eigengap, or the difference between the $i^{th}$ and $(i+1)^{th}$ eigenvalues, plays a crucial role. According to matrix perturbation theory (Stewart, 1990), a larger eigengap indicates a more stable subspace composed of the selected $k$ eigenvectors. Thus, the number of concepts can be determined by identifying the first extreme value of the eigengap  [^4] In this paper, we simply initialize the number of concepts for each window as 3 to obtain the initial representation $\mathbf{Z}$, and then estimate $k$..
@@ Appendix C Proofs and Optimization
C.1 Optimization of Nonconvex Problem